OpenStudy (anonymous):

The expression (3c)^-2 is equivalent to?

OpenStudy (kymber):

5

OpenStudy (anonymous):

?

OpenStudy (anonymous):

The answer could be 1) -6c^2 2) 1/3c^2 3) 1/9c^2 4) 3/c^2

OpenStudy (kymber):

Yeah, five

OpenStudy (anonymous):

There is no 5

OpenStudy (turingtest):

the rule to apply here is\[\large x^{-a}=\frac1{x^a}\]

OpenStudy (turingtest):

make sure that everything in the parentheses is affected by the exponent

OpenStudy (anonymous):

so would it be 1/3c^2

OpenStudy (turingtest):

careful, you need to square \(everything\) in the parentheses

OpenStudy (turingtest):

\[(xy)^a=x^ay^a\]

OpenStudy (anonymous):

[-6c^2]

OpenStudy (turingtest):

combining both rules, we have this\[(xy)^{-a}=\frac1{(xy)^a}=\frac1{x^ay^a}\]

OpenStudy (turingtest):

so there should be no 6 we "distribute" the exponent to everything in the parentheses

OpenStudy (turingtest):

given that\[(xy)^a=x^ay^a\]what is\[(3x)^2\]?

OpenStudy (anonymous):

\[6\times x squared\]

OpenStudy (turingtest):

3^2=9

OpenStudy (anonymous):

\[1 \over 9c^2\]the answer is

OpenStudy (turingtest):

that seems right

OpenStudy (turingtest):

nice job :)

OpenStudy (anonymous):

thanks :)

OpenStudy (turingtest):

welcome